OpenStudy (anonymous):

Reduction (x^2-s^2)^3 / (x-s)^4 (grateful for help!:))

OpenStudy (phi):

the first thing is remember x^2-s^2 can be factored (memorize this!) into (x-s)(x+s) that means your problem is \[ \frac{( (x-s)(x+s))^3}{(x-s)^4} \] can you finish?

OpenStudy (anonymous):

in ^2?? i get it as: ((x-s)(x+s))^3/(x-s)^4 'cause.: (ab)^n = a^n * b^n it becomes: ((x-s)(x+s))^3 = (x-s)^3 * (x+s)^3 (x-s)^3 * (x+s)^3 / (x-s)^4 n > m = ( x + s ) ^ 3 / ( x - s ) ^ ( 4 - 3 ) = (x+s)^3/(x-s) or am I totally wrong?

OpenStudy (phi):

you are totally correct

OpenStudy (phi):

btw, your rule n>m comes from this. if you had \[\frac{x^3}{x^4} = \frac{x\cdot x\cdot x}{x\cdot x\cdot x \cdot x}\] of course you can write that as \[ \frac{x}{x} \cdot \frac{x}{x} \cdot \frac{x}{x} \cdot \frac{1}{x} \] \( \frac{x}{x} \) is 1, so that all simplifies to \[ 1\cdot 1\cdot 1 \cdot \frac{1}{x}= \frac{1}{x} \]

OpenStudy (anonymous):

alright, i see! thank you for your help =D